OpenStudy (anonymous):

Which is equal to ? A. 10−36 B. 10−4 C. 109 D. 1015

OpenStudy (anonymous):

\[\frac{ 10^{12} }{ 10^{-3} }\]

OpenStudy (anonymous):

@Michele_Laino plz help

OpenStudy (anonymous):

hi plz help

OpenStudy (igreen):

Subtract the exponents, and keep the base.

OpenStudy (anonymous):

plz explain too

OpenStudy (anonymous):

don't just give me the answer plz explain

OpenStudy (igreen):

@wilston123 Don't give answers. We are here to guide them to the answer so that they get it on their own. :)

OpenStudy (anonymous):

i am actually start to agree with @iGreen

OpenStudy (igreen):

\(\dfrac{10^{12}}{10^{-3}} \rightarrow 10^{12 - (-3)}\)

OpenStudy (igreen):

Lol.

OpenStudy (anonymous):

ok 12- (-3)=15

OpenStudy (igreen):

Yep, so what's your final answer going to be?

OpenStudy (dmndlife24):



OpenStudy (anonymous):

D? that what i believe lol cool pic

OpenStudy (igreen):

Yep, you got it. \(10^{15}\) is your answer.

OpenStudy (igreen):

Lol, are you talking to me?

OpenStudy (anonymous):